In August 2012, tropical storm Isaac formed in the Caribbean and was headed for the Gulf of Mexico. There was an initial probability of .69 that Isaac would become a hurricane by the time it reached the Gulf of Mexico.

a. What was the probability that Isaac would not become a hurricane but remain a tropical storm when it reached the Gulf of Mexico?

b.Two days later, the National Hurricane Center projected the path of Isaac would pass directly over Cuba before reaching the Gulf of Mexico. How did passing over Cuba alter the probability that Isaac would become a hurricane by the time it reached the Gulf of Mexico? Use the following probabilities to answer this question. Hurricanes that reach the Gulf of Mexico have a .08 probability of having passed over Cuba. Tropical storms that reach the Gulf of Mexico have a .20 probabilit
Answer #1

Answer:---- Date:----12/09/2018

a. Probability that Isaac would not become a hurricane but remain a tropical storm when it reached the Gulf of Mexico 1- 0.69 0.31 b. Given that, P(H) 0.69 P(T) 0.31 P(CIH) 0.08 P(CIT) 0.20 PlHIC) = P(H)P(CIH) / [P(HJP(CIH) + P(T)P(CIT)] = 0.69O.08 / [0.69*0.08 + 0.31*0.20] = 0.4710